OpenStudy (anonymous):

Transform the equation by completing the square: x^2+y^2+6x-16y+48 is greater than or equal to 0. Explanation also please!

OpenStudy (anonymous):

\[(x^2+6x)+(y^2-16y) \ge -48\] \[(x^2+6x+9)+(y^2-16y+64) \ge -48+64+9\] \[(x+3)^2+(y-8)^2 \ge 73-48\] \[(x+3)^2+(y-8)^2 \ge 25\]

OpenStudy (anonymous):

this means you will graph (x+3)^2+(y-8)^2=5^2 and shade everything outside the circle

OpenStudy (amistre64):

completeing a square can be seen as a geometric interpretation; for example: given: x^2 + 6x + ____ , we can "complete a square" like this \begin{array}c &&&(6/2)x\\ &x^2&x&x&x\\ &x&&&\\ (6/2)x&x&&&\\ &x&&&\\ \end{array} now fill in the missing parts to complete it: \begin{array}c &&&(6/2)x\\ &x^2&x&x&x\\ &x&1&2&3\\ (6/2)x&x&4&5&6\\ &x&7&8&9\\ \end{array} 9 missing parts soo... x^2 + 6x +9 is a "complete square

OpenStudy (anonymous):

oh no radar you add 9 on one side you add it to the other just like 64

OpenStudy (amistre64):

since the geometry can get messy, we can stick to the algebra; ax^2 + bx + _____ we take half the x parts (that 'b') and square them to "complete the square" ax^2 + bx + (b/2)^2

OpenStudy (amistre64):

..r^2 = 25 :)